### v2

# Learning 3D-Gaussian Simulators from RGB Videos

Abstract
Realistic simulation is critical for applications ranging from robotics to animation. Learned simulators have emerged as a possibility to capture real world physics directly from video data, but very often require privileged information such as depth information, particle tracks and hand-engineered features to maintain spatial and temporal consistency. These strong inductive biases or ground truth 3D information help in domains where data is sparse but limit scalability and generalization in data rich regimes. To overcome the key limitations, we propose 3DGSim, a learned 3D simulator that directly learns physical interactions from multi-view RGB videos. 3DGSim unifies 3D scene reconstruction, particle dynamics prediction and video synthesis into an end-to-end trained framework. It adopts MVSplat to learn a latent particle-based representation of 3D scenes, a Point Transformer for particle dynamics, a Temporal Merging module for consistent temporal aggregation and Gaussian Splatting to produce novel view renderings. By jointly training inverse rendering and dynamics forecasting, 3DGSim embeds the physical properties into point-wise latent features. This enables the model to capture diverse physical behaviors, from rigid to elastic, cloth-like dynamics, and boundary conditions (e.g. fixed cloth corner), along with realistic lighting effects that also generalize to unseen multibody interactions and novel scene edits. For more information, visit: https://mikel-zhobro.github.io/3dgsim/.
1 Introduction
Figure 1: 3DGSim works directly on multi-view RGB videos and is trained end-to-end on next image prediction. The dynamics model (transformer) operates on particles with latent feature. A learned mapping transforms them into 3D Gaussian Splats for novel view rendering.
Simulating visually and physically realistic environments is a cornerstone for embodied intelligence. Robots must soon tackle tasks like opening washing machines, folding laundry, or tending plants. Traditional analytical simulators demand exact geometry, poses, and material parameters, making arbitrary scene simulation impractical. An alternative is to learn models that predict future states of a scene in large scale observations, as evidenced by the striking visual realism of 2D video generation methods (Li et al., 2022; Wu et al., 2023; NVIDIA et al., 2025). However, pure 2D approaches lack 3D structure awareness, leading to failures in occlusion handling, object permanence, and physical plausibility (Motamed et al., 2025).
3D-based representations address many of these shortcomings, as shown by recent learned simulators which model a wide range of physical phenomena, from fluids and soft materials to articulated and rigid body dynamics Li et al. (2019); Allen et al. (2023). Yet, scaling such methods to data-rich regimes remains challenging as most methods require privileged signals (object-level tracks, depth sensors, physics prior) or hand-crafted graph constructions.
To bridge this gap, we identify three pillars for generalizable, scalable visuo-physical simulation from videos: (1) 3D visuo-physical reconstruction from raw RGB observations; (2) Minimal-bias dynamics modeling to capture diverse physics without explicit biases; (3) Efficient, differentiable decoding back to image space for supervision via reconstruction loss from videos.
Graph neural networks (GNNs) (Sanchez-Gonzalez et al., 2020; Xue et al., 2023; Whitney et al., 2023, 2024; Shi et al., 2024; Wang et al., 2024) have shown great promise in introducing relational inductive biases to handle the unstructured nature of particle sets. This has allowed GNN-based particle simulators to make major progress on all three pillars. In particular, Whitney et al. (2023) jointly train an encoder and dynamics model to learn visuo-physical pixel features from RGBD, and in follow-up work (Whitney et al., 2024) eliminate point correspondences via abstract temporal nodes or per-step models with merging. Driess et al. (2023) demonstrate end-to-end dynamics training of composable NeRF fields from raw RGB images. These advances, in combination with recent advances in feed-forward inverse rendering Chen et al. (2024) and fast differentiable rendering of particles Kerbl et al. (2023), encourage us to ask the question: can we give up the inductive bias arising from locally connected graphs and still learn 3D particle-based simulators?
To this end, we build 3DGSim, a fully end-to-end differentiable framework that embraces the power of scalable computation over hand-crafted biases. 3DGSim begins by inferring 3D visuo-physical features from raw multi-view RGB images through a feed-forward inverse renderer based on MVSplat, extended with a feature extraction module. We then introduce a transformer-only dynamics engine, avoiding kNN-based graph construction and manually designed edge features in favor of learned spatiotemporal embeddings. Finally, a Gaussian Splatting head enables training on an image reconstruction loss from multi-view videos.
Specifically, 3DGSim introduces the following key contributions:
  • Inverse Renderer: Extends MVSplat with a feature extraction module fusing pixel-aligned features into a particle visuo-physical latent representation.
  • Temporal Encoding & Merging Layer: Discards abstract temporal nodes in favor of a hierarchical module that processes an arbitrary number of timesteps.
  • Transformer-Only Dynamics Engine: Removes graph biases and instead uses space-filling curves and learned embeddings for particle-based simulation.
  • End-to-End Differentiable Framework: Connects inverse rendering, transformer dynamics, and Gaussian splatting-based decoding to training for next-frame image reconstruction.
  • Open Source Release: We release the code and dataset to establish a reproducible baseline for future visuo-physical simulation research.
2 Related work
Table 1: Overview on recently proposed particle-based simulators. While most works resort to a combination of kNN and GNNs, our work distinguishes itself by resorting to 3D Gaussian Splatting, space filling curves (SFC) for point cloud serialization, and training the inverse rendering encoder alongside a dynamics transformer.
{NiceTabular}
[…]
3DGSim (Ours) Gaussian splats MVSplat SFC Transformer 3DGS
Encoding and rendering scene representations
Common scene representations include point clouds (particles), meshes, signed distance functions (SDFs), neural radiance fields (NeRFs) Mildenhall et al. (2021), and 3D Gaussians (splats) Kerbl et al. (2023). Point clouds, which approximate object surfaces, can be obtained from RGB-D sensors Shi et al. (2024); Whitney et al. (2023, 2024) or via inverse rendering Wang et al. (2025); Murai et al. (2024); Chen et al. (2024). Several works Whitney et al. (2023, 2024) use U-Net–style encoders trained jointly with the dynamics model, allowing the extracted features to be optimized for physical prediction, a strategy shown to outperform independently trained encoders Li et al. (2022). We adopt this joint training approach using MVSplat Chen et al. (2024), where the encoded features are initially bound to camera parameters. To unify these visuo-physical latents in a global frame, we introduce a learned feature transformation module that maps them into a consistent 3D representation. While many PBS methods render from NeRFs Xue et al. (2023); Shi et al. (2024); Whitney et al. (2023, 2024); Wang et al. (2024); Driess et al. (2023), we instead encode visual appearance directly in the particle cloud using 3D Gaussians. This explicit representation offers high rendering fidelity and significantly improved efficiency over NeRF-based rendering Kerbl et al. (2023), supporting scalability.
GNN based particle-based simulators (PBS)
Graph neural networks (GNNs) introduce relational inductive biases well-suited for modeling the unstructured nature of particle systems. Early work Sanchez-Gonzalez et al. (2020); Li et al. (2019) demonstrated that GNN-based PBS can fit trajectories across a range of physical phenomena. However, GNNs struggle with rigid bodies, where instantaneous velocity changes require long-range message passing across the entire graph in a single step. To address this, later works incorporate mesh structures Pfaff et al. (2021); Allen et al. (2022b) or signed distance functions (SDFs) Rubanova et al. (2024) to enforce object-level coherence. While effective in rigid-body settings, these methods do not generalize easily to deformable or fluid systems. Recent works Saleh et al. (2024); Whitney et al. (2024) suggest adding attention layers to efficiently pass information through the graph. Wang et al. (2024) move toward greater data efficiency by incorporating physics-inspired biases such as the Material Point Method, though limiting broad applicability and requiring small time deltas to simulate. To address temporal correspondence, Whitney et al. (2023) introduces abstract temporal nodes, while Whitney et al. (2024) combines GNNs with transformers to improve memory efficiency by processing and merging pairs of timesteps. However, the method is restricted to two-step horizons, as extending it requires training a separate model for each additional timestep. Typically, these GNN based methods rely on kNN within a fixed radius and hand-crafted features based on object associations and distances to define graph edges and edge features. Message passing and spatial pooling via furthest-point-sampling (FPS) are then used to aggregate information for dynamics prediction. However, kNN and distance computations are expensive and take up 54% of the forward time Wu et al. (2024a), which limits scalability and prevents real-time forecasting. In contrast, we follow the design of PTv3 Wu et al. (2024a). In 3DGSim, we trade off exact KNN neighborhood computation with space-filling curve–based ordering of particles and use sparse convolutions to encode relative positions, avoiding distance calculations. To scale temporally, we propose Temporal Merging with Grid Pooling to construct a hierarchical spatiotemporal, UNet-style Point Transformer for dynamics prediction, abandoning hand-crafted physics in favor of learned representations.
Analytical particle simulators as physical prior
Our work differs in purpose from applications which use Gaussian Splatting particles and analytical PBS as physical prior (e.g. of the shelf differentiable MPM simulator) to accomplish a series of tasks such as tracking Luiten et al. (2024); Keetha et al. (2024); Zhang et al. (2024a); Abou-Chakra et al. (2024), dynamic scene reconstruction Wu et al. (2024b); Huang et al. (2023); Yu et al. (2023) or animation Xie et al. (2023); Zhang et al. (2024b); Lin et al. (2025). While those models can be extended to do parameter estimation via the differentiable nature of simulators Abou-Chakra et al. (2024) they do not fit the generalization of our problem setting.
3 Preliminaries
3DGSim is build atop several prior works, namely: 3D-Gaussian splatting which enables fast rendering, MVSplat which yields 3D Gaussian point clouds from multi-view images, and PTv3 which enables efficient neural processing of 3D point clouds.
Gaussian Splatting
3D Gaussian splatting (3DGS) Kerbl et al. (2023) is an effective framework for multi-view 3D image reconstruction, representation, and fast image rendering and gained rapid popularity due to its support for rapid inference, high fidelity, and editability of scenes. Gaussian splatting uses a collection of 3D Gaussian primitives, each parameterized by
$g_{i}=(p_{i},c_{i},r_{i},s_{i},\sigma_{i}),$ (1)
with the Gaussian’s mean $p_{i}$ (particle position), its rotation $r_{i}$, spherical harmonics $c_{i}$ (defines coloring), scale $s_{i}$, and opacity $\sigma_{i}$. To render novel views, these primitives are projected onto a 2D image plane using differential tile-based rasterization. The color value at pixel $\mathbf{p}$ is calculated via alpha-blend rendering: $I(\mathbf{p})=\sum_{i=1}^{N}\alpha_{i}c_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ where $\alpha_{i}=\sigma_{i}e^{-\frac{1}{2}(\mathbf{p}-p_{i})^{{}^{\top}}\Sigma_{i}^{-1}(\mathbf{p}-p_{i})}$ is the 2D density, $I$ is the rendered image, $N$ is the number of primitives in the image, and $\Sigma_{i}$ is the covariance matrix given by $\Sigma_{i}=r_{i}s_{i}r_{i}^{{}^{\top}}$ for improved computational stability.
Figure 2: MVSplat uses a cost volume with plane sweeping to regress pixel-wise 3D Gaussians, which are unprojected to world frame using camera parameters.
MVSplat: Multi-view images to 3D point clouds
MVSplat deploys a feed-forward network $f_{\phi}$ with parameters $\phi$ that maps $M$ images $\mathcal{I}=\{I^{m}\}_{i=m}^{M}$ with $I^{m}\in\mathbb{R}^{(H\times W\times 3)}$ to a set of 3D Gaussian primitives (Fig.˜2)
$f_{\phi}:\{(I^{m},P^{m})\}_{m=1}^{M}\mapsto\{g_{i}\}_{i=1}^{M\times H\times W}.$
At each time step, MVSplat localizes Gaussian centers using a cost volume representation through plane sweeping and cross-view feature similarities. To do so, it requires the corresponding camera projection matrices $\mathcal{P}=\{P^{m}\}_{m=1}^{M}$ which are calculated as $P^{m}=K^{m}[R^{m}|t^{m}]$ via the matrices for the camera intrinsics and extrinsics $K^{m}$, rotation $R^{m}$, and translation $t^{m}$.
4 3DGSim
3DGSim is a fully differentiable pipeline that, given $T$ past multi-view RGB frames, reconstructs 3D particles with latent features, simulates their motion, and renders the next frames. It consists of three jointly trained modules (Fig.1): (i) an encoder that maps multi-view RGB images to 3D particles, (ii) a dynamics model that simulates the motion of these particles through time, and (iii) a renderer that yields images by first mapping the particles to Gaussian splats.
4.1 State Representation
To simulate physical scenes from vision, we require a state representation that is both expressive enough to capture fine-grained 3D and physical properties, and compact enough to enable efficient learning and prediction. While an explicit 3DGS representation $g_{i}(t_{k})$ offers geometric and visual completeness, it is insufficient for dynamics modeling. Instead, we distill the state of each particle into a more compact representation:
$\tilde{g}_{i}(t_{k})=\bigl{(}p_{i}(t_{k}),\,f_{i}(t_{k})\bigr{)}$ (2)
where $t_{k}$ denotes the k-th timestep and $f_{i}\in\mathbb{R}^{d}$ the visuo-physical latent particle feature, encoding shape, appearance, and dynamic properties. Unless otherwise stated, we omit the timestep $t_{k}$ and the particle index $i$ when the statement applies to all timesteps or particles, respectively.
Optional: Masking and Freezing of Particles
At each timestep $t_{k}$, the encoder yields pixel-aligned features for each input image. As an optional step, we apply a foreground mask to discard particles likely belonging to the static background, retaining a reduced set of $N_{k}$ particles per time step (Fig. 2). Additionally, as originally suggested by Whitney et al. (2023), static particles can optionally be “frozen”, i.e. act as input to the dynamics model but are excluded from position updates. These optional strategies improve efficiency without being necessary for successful training, as shown in Sections˜5 and B.
Figure 3: Position $p$ and dynamic features $f^{\text{dyn}}$ are updated while $f^{\text{inv}}$ remain constant.
Invariant and dynamic feature decomposition
We decompose each particle’s visuo-physical feature into an invariant and a dynamic part as shown in Fig.˜3, writing
$f_{i}=f_{i}^{\text{inv}}\oplus f_{i}^{\text{dyn}},$
where $\oplus$ denotes concatenation. The dynamics model updates only $f_{i}^{\text{dyn}}$, while leaving $f_{i}^{\text{inv}}$ unchanged. For clarity, we will refer to the dynamics update as “updating $f_{i}$”, though only the dynamic component $f_{i}^{\text{dyn}}$ is altered.
4.2 View-Independent Inverse Renderer
In the proposed augmentation of MVSplat, the pixel-aligned features $\hat{f}_{i}^{\prime}$ are tied to the view-dependent camera frame in which they were observed. This camera-centric encoding hinders generalization and leads to subpar model performance as dynamics prediction should ideally be invariant to the observer’s viewpoint. To address this, 3DGSim extends MVSplat with a learned feature encoding network that transforms pixel-aligned features $\hat{f}_{i}^{\prime}$ into view-independent latent features $f_{i}$. The feature encoding network leverages the pixel depth, pixel shift, density, and ray geometry, represented via Plücker coordinates Plücker (1868-1869) to infer spatially consistent 3D features via FiLM conditioning Perez et al. (2017). This yields an inverse rendering encoder that outputs canonically anchored particle states suitable for downstream dynamics learning. Details are provided in Section˜A.1.
4.3 Dynamics Model
At the core of our method is the dynamics model, a transformer architecture operating on particle sets in space and time. The dynamics model receives as input the full particle trajectory
$\big{\{}\{\tilde{g}_{i}(t_{k})\}_{i=1}^{N_{k}}\big{\}}_{k=1}^{T},\quad\text{where }\tilde{g}_{i}(t_{k})=\big{(}p_{i}(t_{k}),f_{i}^{\text{inv}}(t_{k}),f_{i}^{\text{dyn}}(t_{k})\big{)},$ (3)
and predicts the updated dynamic features at the next timestep
$\Delta p(t_{T}),\Delta f^{\text{dyn}}(t_{T})=\mathrm{Dynamics\,Model}\left(\{\{\tilde{g}_{i}(t_{k})\}_{i=1}^{N_{k}}\}_{k=1}^{T}\right),$ (4)
such that $p_{i}(t_{T+1})=p_{i}(t_{T})+\Delta p_{i}(t_{T})$ and $f_{i}^{\text{dyn}}(t_{T+1})=f_{i}^{\text{dyn}}(t_{T})+\Delta f_{i}^{\text{dyn}}(t_{T})$. As these point clouds are unstructured and potentially vary in size at each time step due to masking, a fundamental challenge arises: How can a network efficiently propagate the embedded physics information both spatially and temporally?
We tackle this question by building on PTv3 Wu et al. (2024a), which has recently achieved state-of-the-art performance in representation learning for unstructured point clouds. As discussed in Section˜A.2, PTv3 operates by serializing the input point cloud and applying patch-wise attention. However, the original design of PTv3 is limited to point clouds that do not exhibit temporal variation. In this section, we extend PTv3 to predict dynamics from temporally evolving point clouds. First, we extend serialization to equip point cloud encodings with a timestamp. Then, we equip features with temporal embeddings that allow attention to distinguish timestamps. Lastly, we use the timestamps to merge temporally adjacent features that can then be processed by PTv3’s patch-wise attention blocks.
Figure 4: The dynamics model encodes the time step into each embedding and merges embeddings from adjacent timesteps. The TEM and PTv3 blocks are applied repeatedly until all embeddings are merged. Our extensions to PTv3 are highlighted in red.
Figure 5: Temporal point cloud serialization.
Figure 6: Temporal merging and embedding followed by patch grouping for applying patch-wise attention.
Temporally serialized point cloud (t-SPC)
To enable spatio-temporal reasoning over multiple timesteps, we extend PTv3’s point serialization scheme by encoding both spatial and temporal structure into a single key. Specifically, for each particle $i$ at timestep $t_{k}$ in batch $b$, we define a 64-bit serialization code:
$\tilde{s}_{i}(t_{k},b)=\big{[}\underbrace{b}_{\text{(64 - $\tau$ - $\kappa$)\,Bits}}\,|\,\underbrace{s_{t_{k}}}_{\text{$\tau$\,Bits}}\,|\,\underbrace{s_{i}}_{\text{$\kappa$\,Bits}}\big{]}$ (5)
Here, $s_{t_{k}}$ is the temporal code and $s_{i}$ is a spatial code being obtained by projecting $p_{i}$ onto a space-filling curve (SFC). We set $\kappa=48$ and allocate $\tau=\log_{2}(T)$ bits for time. With 16 bits per dimension and a grid resolution of $G=0.004\,\mathrm{m}$, the spatial encoding spans up to $216\,\mathrm{m}$ per axis.
Temporal encoding
As shown in Figure 6, before merging t-SPCs across timesteps, we inject a learned, timestep-specific positional encoding $E_{t_{k}}$ as
$f_{i}(t_{k})\leftarrow f_{i}(t_{k})+E_{t_{k}}.$ (6)
This temporal encoding ensures that the attention mechanism can distinguish points across different temporal instances, enabling the model to reason about dynamics over time. Similar positional encoding methods have previously been applied in transformer architectures to differentiate positions within sequences Vaswani et al. (2017).
Temporal merging
Unlike PTv3, which limits attention exclusively to patches composed of points from the same time step, our method extends the receptive field across time steps. To do so, we propose temporal merging, which applies a one-bit right shift to the temporal codes $s_{t_{k}}$:
$\mathrm{Merge}(\tilde{s}_{i})=[b\mid(s_{t_{k}}\gg 1)\mid s_{i}].$ (7)
For example, consider points initially serialized with temporal codes $s_{t_{1}}=0$ and $s_{t_{2}}=1$ as depicted in Figure 6. After temporal merging, the updated temporal codes become $s_{t_{1}}=0$ and $s_{t_{2}}=0$. Thus, the points originally belonging to separate time steps now receive identical batch-temporal codes, enabling patch grouping to form patches that contain points from multiple time steps simultaneously. Consequently, our attention module can effectively model relationships among points across adjacent time steps within a single computational step.
In our current implementation, we merge pairs of consecutive time steps $t_{j}$ and $t_{j+1}$, where $j\in 2n,n\in\mathbb{N}^{*}$. We adopted this simple and computationally convenient strategy, though more elaborate temporal merging schemes – such as aggregating points from larger windows of adjacent time steps – can also be explored in future work. Importantly, while Whitney et al. (2024) deploy a dedicated transformer module for each time step, our proposition of temporal merging enables the reuse of the same attention block across time steps. This design significantly reduces memory consumption while also promoting knowledge transfer between time steps.
Patch-wise attention and particle-wise MLP
After each temporal encoding and merging (TEM) block, the cloud embeddings are processed by PTv3’s patch-wise attention block. First, the embeddings are equipped with a position encoding via a sparseCNN with skip connection (xCPE in Fig.˜4). Then, the embeddings are fed to a patch-wise attention layer. Finally, at the end of the dynamics model, each particle alongside its embedding is mapped by a particle-wise MLP to $\Delta p_{T}$ and $\Delta f_{T}$.
4.4 Rendering Features for the Image Reconstruction Loss
To render images with 3DGS, particle states $\tilde{g}_{i}=(p_{i},f_{i})$ are transformed into Gaussian splat parameters $g_{i}$ via a learned head, materialized only at the final stage to supervise the training with image reconstruction.
3DGSim is trained solely on an image reconstruction loss $\mathcal{L}$. This loss is computed from rasterized multi-view images, generated based on both the encoder predictions of past point clouds $\{\{g_{i}(t_{k})\}_{i=1}^{N_{k}}\}_{k=0}^{T}$ and the simulated future point cloud trajectory $\{\{g_{i}(t_{k})\}_{i=1}^{N_{k}}\}_{k=T+1}^{T+T^{\prime}}$. Specifically, the loss reads
$\mathcal{L}=(1-\lambda)\frac{1}{T}\sum_{k=0}^{T}\mathcal{L}_{k}+\lambda\sum_{k=T+1}^{T+T^{\prime}}\gamma^{k-T-1}\mathcal{L}_{k}\quad\text{and}\quad\mathcal{L}_{k}=\mathcal{L}_{2}(I_{k}^{\text{gt}},I_{k})+\beta\,\mathcal{L}_{\text{LPIPS}}(I_{k}^{\text{gt}},I_{k}),$ (8)
with $\lambda=0.5$, temporal decay factor $\gamma=0.87$, $T\in\{2,4\}$ and $T^{\prime}=12$. The per-frame reconstruction loss $\mathcal{L}_{k}$ measures the discrepancy between ground-truth ($I_{k}^{\text{gt}}$) and predicted ($I_{k}$) multi-view images using a weighted combination of pixel-wise $\ell_{2}$ and LPIPS Zhang et al. (2018) terms with hyper-parameter $\beta=0.05$.
Figure 7: Qualitative examples of 3DGSim’s dynamic predictions. The examples show accurate modeling of deformations, rigid motion, cloth with corner constraints, and shadows, despite training on less than 6 minutes of video per object across 6 objects.
Figure 8: Trajectory PSNR of 3DGSim and Cosmos. The novel view reconstruction PSNR is shown for both past and future predictions. The Cosmos model is conditioned on 9 past frames and appropriate language prompts.
Figure 9: 3DGSim’s prediction of a rigid plane captures shadows by altering ground particle appearance.
Figure 10: Although 3DGSim has not been trained on scenes containing this particular elastic object nor several objects, it predicts physically plausible deformations. Interestingly, the combined weight of five objects slowly pushes the particles through the table causing color changes or falling particles.
Figure 11: Comparison of 3DGSim to Cosmos. Explicit models use 3DGS parameters and a static latent feature as inputs to the dynamics model, while latent models use only latent features mapped to Gaussians after dynamics. The $\dagger$-model uses only 6 camera views (3 input + 3 reconstruction) instead of 12 (4+5); the $\ddagger$-model omits segmentation masks for static elements. Metrics future and past are means over all timesteps. "4-12" means 4 past steps predicting 12 future steps. CosmosFT refers to LoRA fine-tuned version of the Cosmos-Predict2 2B model on the respective dataset.
{NiceTabular}
llcccc Dataset Model PSNR (future) $\uparrow$ PSNR (past) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
[…]
3DGSim 2-12 latent 28.08 ± 2.46 33.00 ± 1.62 0.90 ± 0.03 0.09 ± 0.03
3DGSim 4-12 explicit 27.88 ± 2.43 32.77 ± 1.57 0.90 ± 0.03 0.09 ± 0.03
3DGSim 2-12 explicit 27.07 ± 2.27 32.67 ± 1.65 0.90 ± 0.03 0.09 ± 0.03
CosmosFT 26.44 ± 2.26 – 0.68 ± 0.05 0.10 ± 0.03
Cosmos 22.35 ± 3.82 – 0.83 ± 0.08 0.24 ± 0.08
[…]
3DGSim 2-12 latent 32.05 ± 3.48 35.99 ± 1.88 0.96 ± 0.02 0.03 ± 0.02
3DGSim 2-12 explicit 29.92 ± 1.72 40.85 ± 2.94 0.96 ± 0.02 0.03 ± 0.02
3DGSim 4-12 explicit 29.69 ± 1.75 40.16 ± 3.07 0.97 ± 0.02 0.02 ± 0.01
3DGSim 4-12 latent $\dagger$ 31.60 ± 3.09 32.55 ± 2.12 0.97 ± 0.02 0.02 ± 0.01
3DGSim 4-12 latent $\ddagger$ 32.66 ± 3.43 34.45 ± 2.44 0.96 ± 0.02 0.03 ± 0.02
CosmosFT 26.50 ± 5.21 – 0.82 ± 0.02 0.07 ± 0.03
Cosmos 18.87 ± 3.99 – 0.79 ± 0.08 0.23 ± 0.08
[…]
3DGSim 2-8 latent 26.25 ± 2.38 35.22 ± 1.97 0.88 ± 0.03 0.08 ± 0.02
3DGSim 4-8 explicit 23.72 ± 1.52 39.75 ± 2.32 0.89 ± 0.03 0.08 ± 0.03
3DGSim 2-8 explicit 17.97 ± 2.02 35.47 ± 1.68 0.88 ± 0.03 0.08 ± 0.02
CosmosFT 22.49 ± 0.99 – 0.73 ± 0.03 0.14 ± 0.04
Cosmos 21.10 ± 3.56 – 0.86 ± 0.06 0.19 ± 0.06
5 Experiments
In what follows, we train 3DGSim on different datasets and test the model’s ability to generalize.
Model setup
Unless stated otherwise, the following training and parameter settings serve as defaults in the experiments. The state consists of dynamic $f^{\text{dyn}}$ and invariant features $f^{\text{inv}}$ of size $(32,32)$ for the implicit- and $(n_{f},16)$ for the explicit 3D Gaussian particle representation. In the explicit representation, $f^{\text{dyn}}$ corresponds to explicit Gaussian primitives of size $n_{f}$ which are directly used for rendering. The inverse rendering encoder follows MVSplat, reducing candidate depths from 128 to 64 due to smaller scene distances. Default near-far depth ranges are [0.2, 4] for rigid bodies and [1.5, 8] for the other datasets, as the scene has a larger scale. The dynamics transformer defaults to PTv3 with a 5-stage encoder (block depths [2,2,2,6,2]) and a 4-stage decoder ([2,2,2,2]). Grid pooling and temporal merging strides default to [1,4,2,2,2] and [1,2,2,2,2], respectively, with grid size G=0.004G{=}0.004 m. Attention blocks use patches of size 1024, encoder feature dimensions [32,64,128,256,512], decoder dimensions [64,128,256], encoder heads [2,4,8,16,32], and decoder heads [4,4,8,16]. For the camera setup, we select 4 uniformly distributed views at random and an additional 5 target cameras from the remaining cameras (out of 12 total) to compute the reconstruction loss.
Training
Our models are trained with AdamW for $\sim$120,000 steps using a cosine annealing warm-up and a learning rate of $2\times 10^{-4}$, with batch sizes of 6 and 4 for 2-step and 4-step states, respectively. To optimize memory and speed, we use gradient checkpointing and flash attention v2 Dao (2024). Training is performed on a single H100 GPU and typically takes around 5 days..
Datasets
To evaluate 3DGSim’s robustness in learning dynamics from videos, we introduce three challenging datasets: rigid body, elastic, and cloth.
The rigid body dataset consists of 1,000 simulated trajectories involving six rigid objects (turtle, sonny school bus, squirrel, basket, lacing sheep, and turboprop airplane) from the GSO dataset Downs et al. (2022). Each trajectory spans 32 frames at 12 FPS, providing controlled dynamics characteristic of rigid body motion. The elastic dataset, aimed at capturing plastic deformable object dynamics, includes six objects (dragon, duck, kawaii demon, pig, spot, and worm) simulated using the Genesis MPM elastoplastic simulator Authors (2024). Each object undergoes deformation upon collision with a circular gray ground, offering scenarios of complex elastic behavior. The cloth dataset includes the same set of objects as the elastic dataset. Here, the cloth is fixed at four corners, posing the challenge to infer implicit constraints and modeling dynamic cloth-like deformations.
Both elastic and cloth datasets include 200 trajectories per object, simulated with a 0.001 time step and 20 substeps. Each two second sequence is recorded at 42 FPS resulting in 84 frames per trajectory and less than 6 minutes of footage per object.
Benchmarking
Existing baselines do not allow direct comparison without substantial reimplementation. Key methods – VPD, HD-VPD, DEL, and 3D-IntPhys Whitney et al. (2023, 2024); Wang et al. (2024); Xue et al. (2023)lack public code and data, also unavailable upon contacting authors. Without published datasets, any reimplementation would lack verifiability, limiting reproducibility and fair evaluation. To address this, we will release our code and datasets. DPI-Net and VGPLDP Li et al. (2019, 2020) are open-source but rely on ground-truth particle trajectories and require major adaptation to fit our setting. Thus, we provide quantitative comparison to Cosmos NVIDIA et al. (2025), noting it’s not directly comparable to our multi-view setup as this model is pretrained while receiving multiple past frames from a single view. For fair comparison, we evaluated both the base model and a LoRA-finetuned variant (CosmosFT) trained for 6,000 iterations using recommended parameters. Cosmos is conditioned on prompts: “A rigid body falling on a circular gray ground”, “A soft body falling on a circular-gray-ground”, and “A rigid body falling on a red rectangular cloth which is fixed on its corners”.
Metrics
For evaluation, 12% of trajectories are chosen at random and held out from each dataset, and we report each model’s PSNR, LPIPS and SSIM.
5.1 Trajectory Simulation
Figure 8 shows the PSNR of our model’s simulations using two or four past frames and the latent representation. To the best of our knowledge, 3DGSim is currently the only open-source, particle-based simulator trained directly on videos that integrates rendering and simulation; therefore, we compare against Cosmos-1.0-Autoregressive-5B-Video2World NVIDIA et al. (2025), a video generation model conditioned on 9-frame video input and language prompts. Additionally, we ablate key hyperparameters and report results in Table 4.4. Explicit representations achieve similar performance but generalize poorly, particularly with fewer cameras (see supplementary).
5.2 Scene Editing and Model Generalization
A key advantage of 3DGSim is its 3D representation of the simulator’s state, enabling direct scene editing for modular construction, counterfactual reasoning, and scenario exploration. By modifying object placements, adding elements, or altering physical properties, we test how the model adapts, demonstrating its flexibility and interpretability. To illustrate the model’s editability, we modified the height of the ground or removed it entirely – scenarios not encountered during training. Remarkably, the model adapts seamlessly, continuing to produce realistic simulations, which suggests a robust understanding of physics beyond its training distribution. For further demonstration, the interested reader is referred to the accompanying video.
As an additional rigorous test, we simulate multi-object interactions by duplicating bodies (see Figure 4.4 and Section˜C.1), and run the simulation for 180 steps. Despite being trained only on object-ground collisions, 3DGSim correctly captures realistic multi-body dynamics. Instead of collapsing into chaotic interactions, individual objects retain structural integrity and move cohesively. We attribute this to the transformer architecture, which avoids structural constraints that limit generalization.
A striking consequence of removing explicit physics biases is that 3DGSim not only captures physics but also learns to reason about broader scene properties. One particularly intriguing outcome is its ability to model shadows as shown in Figure 4.4, which requires a holistic understanding of scene geometry, lighting, and object relationships.
Figure 12: Prediction speed of 3DGSim versus simulation FPS.
5.3 Simulation Speed
Simulation speed is critical for robotics applications. Traditional simulators (FEM, MPM, PBD) typically employ small integration timesteps. Learned approaches enable larger timesteps, allowing 3DGSim to simulate elastic cloth at 42 FPS and rigid dynamics at 12 FPS, with inference speeds of ${\sim}16$ FPS (4 past steps) and ${\sim}20.1$ FPS (2 past steps), using under 20 GB VRAM on an H100 GPU and achieving near real-time prediction speeds as illustrated in Fig.˜12.
6 Discussion
We introduced 3DGSim, a fully differentiable 3D Gaussian simulator that operates directly on multi-view RGB video. Our framework unifies inverse rendering, dynamics prediction, and novel-view video synthesis into an end-to-end learnable system. By extending MVSplat to encode both visual and physical information into particle-centric latent features, 3DGSim bridges perception and simulation. Unlike prior particle-based simulators that depend on GNNs and kNN graph construction, our transformer-based dynamics model leverages scalable computation with space-filling curves and learned temporal embeddings to reason over diverse physical interactions. This enables 3DGSim to generalize on rigid, deformable, and compound object dynamics without object-level supervision.
By avoiding inductive biases where possible and prioritizing scalability over first-principles, transformers have already demonstrated their effectiveness in uncovering complex patterns within large datasets. In turn, we are eager to explore 3DGSim on significantly larger datasets encompassing a diverse range of physical interactions. As we scale up, incorporating language embeddings into simulators emerges as a particularly promising direction. After all, words such as “liquid” or “mirror” could provide meaningful priors about an object’s physical properties. In this regard, we hope that 3DGSim points the way towards scalable particle-based simulators that are able to infer the interactions of complex physical phenomena from video as well as potentially text data, paving the way for robots that truly grasp the dynamics of their surroundings.
Limitations
3DGSim relies on multi-view data, which remains scarce at scale. This limitation could be mitigated by leveraging monocular inverse rendering models such as Wang et al. (2025); Murai et al. (2024). While occlusions are not explicitly handled, the dynamics model partially compensates for missing data; this could be further improved through point completion. Action conditioning, as demonstrated in Whitney et al. (2024), can also be embedded and treated similarly to other particles. Finally, validation on real-world datasets is left for future work as this study primarily serves as a proof of concept.
Appendix A 3DGSim additional details
A.1 Unprojecting Pixel-Aligned Features via FiLM Conditioning
To transform the view-dependent pixel-aligned features $\hat{f}_{i}^{\prime}$ into a consistent 3D representation, we use a multilayer perceptron (MLP) with Feature-wise Linear Modulation (FiLM). FiLM conditioning enables the MLP to adapt its processing of $\hat{f}_{i}^{\prime}$ based on geometric context, such as camera viewpoint and depth.
Specifically, FiLM computes a scale and bias using a conditioning network $\gamma$ that takes as input a geometric context vector $\mathbf{x}_{i}$—which includes depth, density, and pixel shift—as well as the Plücker ray encoding $\mathbf{r}_{i}=[\mathbf{o}_{i}\times\mathbf{d}_{i}\mid\mathbf{d}_{i},]$, where $\mathbf{o}_{i}$ and $\mathbf{d}_{i}$ denote the origin and direction of the viewing ray:
$\text{scale}_{i},\,\text{bias}_{i}=\gamma(\mathbf{x}_{i},\,\mathbf{r}_{i}).$ (S1)
These parameters modulate the activations of the MLP processing $\hat{f}_{i}^{\prime}$ through FiLM layers:
$f_{i}=\text{MLP}(\hat{f}_{i}^{\prime};\,\text{scale}_{i},\,\text{bias}_{i}),\quad\text{FiLM}(h)=\text{scale}_{i}\odot h+\text{bias}_{i},$ (S2)
where $h$ denotes a hidden activation and $\odot$ is element-wise multiplication.
This setup allows the network to unproject image-aligned features into canonical 3D space while respecting scene geometry and view direction.
A.2 PTv3: Scalable point cloud transformations
Point cloud serialization
At the core of PTv3 lies “point cloud serialization”, an algorithm that transforms an unstructured point cloud into ordered points. This process begins by discretizing 3D space into a uniform grid of points. As illustrated in Figure 5, these point are then connected using a space-filling curve – a path that traverses each grid point exactly once while preserving spatial proximity as much as possible. Each point $p_{i}$ is assigned an integer code $s_{i}$, representing its position within a space-filling curve, via the mapping
$s_{i}=\phi^{-1}(\lfloor p_{i}/G\rfloor)$ (S3)
with $\phi^{-1}:\mathcal{Z}^{N}\mapsto\mathcal{Z}$ and grid size $G\in\mathbb{R}$. The points in the clouds are then ordered by their respective code $s_{i}$, yielding a serialized point cloud (SPC)
$S_{i}(t_{k})=\{(s_{i},\tilde{g}_{i})\}_{i=1}^{N_{k}}.$ (S4)
While this approach may not preserve local connectivity as precisely as kNN groupings, Wu et al. (2024a) emphasizes that the slight loss in spatial precision is outweighed by a significant gain in computational efficiency. To obtain diverse spatial connections between points, PTv3 shuffles between four different space filling curve patterns to obtain SPCs from which patches are computed and varies the patch computation through integer dilation.
Patch grouping
PTv3 partitions the SPC $S_{i}(t_{k})$ into equally sized patches and applies self-attention within each patch. To ensure divisibility, patches that do not align with the specified size are padded by borrowing points from neighboring patches.
Conditional embeddings and patch attention
Besides the computational efficiency of SPC over kNN, its main advantage lies in the compatibility with standard dot-product attention mechanisms. To understand this, we first examine how the standard PTv3 architecture computes particle-wise predictions from a single point cloud at a single time step. The process begins by extracting particle-wise embeddings $E_{i}$ for each serialized point $(s_{i},\tilde{g}_{i})$ using a sparse convolutional neural network (CNN). Next, the embedded SPCs are progressively down sampled via grid pooling before being grouped and shuffled into patch pairs. Conditional positional embeddings (xCPE) are then added to the embeddings, followed by layer normalization and a patch-wise attention layer predicting the change in the embeddings $\Delta E_{i}$. The pooling, patch shuffling, and attention blocks are arranged in a U-net Ronneberger et al. (2015) like architecture that first reduces the size of the SPCs in an encoding step and then mirrors this architecture. In 3DGSim, the final layer of the dynamics model predicts the change in the particle positions $\Delta p_{i}$ and the change in their features $\Delta f_{i}$.
This section describes additional implementation details that were not explicitely mentioned in the main paper.
Feature Encoding Network
In 4.2 we describe the feature encoding network that transforms pixel-aligned features into view-independent the latent features. To regress the FiLM conditioning we use a 2-layer CNN with GELU activation, kernel of size 3 and channel dimensions $(10,20)$. The first dimension also matches the size of the conditioning vector.
Particle Wise MLP
At the end of the dynamics model, the embedding of each particle is mapped back to the particle latent features. For that we use an MLP with shapes $(128,128)$ and GELU activation between each layer.
Appendix B Ablations
For the ablation studies of 3DGSim, only the elastic dataset is used. The default configuration uses the latent representation, 4-step past conditioning, 12-step future rollouts, 4 input views, 5 target views, and a total of 12 cameras for training, unless otherwise specified. Any deviations from these parameters are explicitly stated or made clear within the context of the respective ablation.
This analysis provides detailed insights into design and strategic choices that inform future improvements of our approach.
B.1 Rollout Length.
Rollout Length PSNR Future PSNR Past
2 steps 26.43 ± 3.48 35.70 ± 2.51
4 steps 28.83 ± 4.96 35.25 ± 2.53
8 steps 30.64 ± 3.23 33.86 ± 2.17
12 steps 33.15 ± 3.51 34.55 ± 2.26
Table S1: Rollout Length
We evaluate the influence of prediction rollout length during training (2, 4, 8, and 12 steps). Consistent with expectations, results improve significantly as the rollout length increases, reaching a peak performance at 12 steps with a PSNR Future of $33.15\pm 3.51$. Extending the number of rollout steps enhances the model’s predictive capability but leads to significant memory requirements. Employing regularization methods such as random-walk noise injection or diffusion techniques or even other modality (see later) can help reduce the required rollout steps.
B.2 Camera Setup.
Setup PSNR Future PSNR Past
Explicit 3 views out of 6 21.02 ± 1.78 16.86 ± 0.83
Latent 3 views out of 6 31.60 ± 3.09 32.55 ± 2.12
Latent 4 views out of 12 33.15 ± 3.51 34.55 ± 2.26
Table S2: Camera Setup (85k steps)
To approximate a realistic scenario suitable for real-world deployment, we investigate performance with reduced camera setups. Interestingly, the latent representation models achieve robust performance even when trained with 3 views out of 6 total cameras (PSNR 33.15), whereas explicit representation models degrade significantly (PSNR drops to 21.02) due to convergence of the encoder to poor local minima. This local minimum manifests as camera-specific overfitting, where the model erroneously predicts particle arrangements forming planar, screen-aligned shapes. While this artificially reduces the training loss of target viewpoints, it undermines the true representation quality and disrupts convergence to a consistent 3D reconstruction. Further restricting the setup to only 2 views out of 4 cameras results in unsuccessful training for both latent and explicit models, indicating that very limited camera setups demand careful placement or preliminary encoder pre-training, which should be investigated in a future work.
B.3 Segmentation Masks.
Segmentation PSNR Future PSNR Past
Without masks 32.66 ± 3.43 39.08 ± 3.18
With masks 33.15 ± 3.51 34.55 ± 2.26
Table S3: Segmentation Masks
While our final models use segmentation masks for static objects (e.g., ground surfaces), we explore training without these masks to test model reliance on explicit segmentation. We find only a slight reduction in performance (32.66 vs. 33.15), demonstrating the models’ capability to implicitly infer static scene regions directly from raw RGB inputs. Thus, explicit segmentation masks are helpful but not strictly essential.
B.4 Modality Configurations
Modality PSNR Future PSNR Past
4-1-2-6 32.59 ± 3.22 33.39 ± 2.21
4-4-1-3 31.98 ± 3.78 34.03 ± 2.39
4-1-1-12 33.15 ± 3.51 34.55 ± 2.26
Table S4: Input Modality
Different input modality configurations were tested. Here, we adapt the notation "a-b-c-d", each varying the temporal span of input conditioning $a$ is the number of particle frames representing the state, $b$ indicates the number of backward frames used to predict $c$ future steps, and total rollout steps during training $d$, leading to $b\cdot c\cdot d$ rollout steps per training step. Both variants ("4-1-2-6", "4-4-1-3" attain competitive performance (PSNR of 32.59 and 32.27 respectively), significantly reducing the computational load compared to longer standard rollouts. This highlights promising avenues for future investigation, emphasizing balance between computational efficiency and performance quality.
B.5 Grid Resolution
Grid Size PSNR Future PSNR Past
0.002 25.39 ± 2.95 32.28 ± 2.68
0.004 33.15 ± 3.51 34.55 ± 2.26
0.008 25.40 ± 3.84 30.78 ± 2.63
0.0012 24.06 ± 3.53 31.74 ± 2.56
Table S5: Grid Resolution.
We test a series of grid resolutions (0.002, 0.004, and 0.008), observing optimal results at 0.004 with a PSNR of 33.15. Both higher (0.008) and finer resolutions (0.002) degrade performance, suggesting an optimal balance achieved at 0.004 between detail preservation and computational complexity for out scene size.
B.6 Temporal Merger.
Temporal Merger Setup PSNR Future PSNR Past
[1,1,2,2..] with embedding 27.55 ± 3.22 31.68 ± 2.60
[1,1,4,..] with embedding 26.79 ± 2.94 32.21 ± 2.73
[1,2,2,..] without embedding 25.07 ± 3.22 31.16 ± 2.59
[1,2,2,..] with embedding (120k) 33.15 ± 3.51 34.55 ± 2.26
[1,1,1,2,2] with embedding 18.87 ± 1.50 18.09 ± 1.45
[1,4,..] with embedding 18.19 ± 1.29 18.33 ± 1.48
Table S6: Temporal Merger. The models are trained for 80k steps if not specified otherwise.
We experiment with various temporal merging strides for each encoder stage combined with embedding options of timestep position encoding. Since we only train with 4 past steps, the strides ".." don’t influence the results. After 80k iterations, results clearly indicate two critical factors for success: the use of learned positional embeddings and timing of merging operations. Optimal results occur when merging temporal information only after early spatial processing stages ([1,2,2..]), whereas early or too-late merging drastically diminishes performance. Poor outcomes with late merging likely arise due to spatial pooling operations that dilute vital temporal distinctions before merging.
Appendix C Visualizations
In this section we show rolled out dynamic prediction for different scenarios. If not otherwise stated, the first row in each image shows the ground truth predictions. First we depict scene editability and composabilty followed by several generalizations to multi body and visualization for each dataset. Videos of the predictions below are available in the supplementary material.
C.1 Scene editability
Figure S1: Ground Removal
Figure S2: Higher Throw
Figure S3: Parallel Simulation
C.2 Generalization to multiple bodies
Figure S4: Scene with 2 bunnies.
Figure S5: Scene with 3 bunnies.
Figure S6: Scene with 5 bunnies.
Figure S7: Scene with 5 worms.
Figure S8: Various simulation scenarios with different numbers and types of objects.
C.3 Visualization of cloth simulations
Figure S9: Dragon
Figure S10: Duck
Figure S11: Cow
Figure S12: Devil
Figure S13: Various objects falling on a red cloth with its corners fixed.
C.4 Visualization for elastic dynamics
Figure S14: Cow
Figure S15: Dragon
Figure S16: Duck
Figure S17: Pig
Figure S18: Worm
Figure S19: Various elastic simulations for different objects.
C.5 Visualization for rigid body dynamics
Figure S20: Car
Figure S21: Squirrel
Figure S22: Shoe
Figure S23: Box
Figure S24: Turtle
Figure S25: Various rigid simulations for different objects.